OpenStudy (nuccioreggie):

Simplify negative 3 over 5 ÷ 7 over 6. negative 7 over 10 negative 18 over 35 18 over 35 7 over 10

OpenStudy (igreen):

\(-\dfrac{3}{5} \div \dfrac{7}{6}\) When dividing fractions, we flip the 2nd fraction and multiply. \(-\dfrac{3}{5} \times \dfrac{6}{7} \rightarrow \dfrac{-3 \times 6}{5 \times 7}\) Can you simplify that?

OpenStudy (nuccioreggie):

-18 n 35

OpenStudy (igreen):

Yes, so we have -18/35

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

plz

OpenStudy (igreen):

@IQ2121 Make your own question

OpenStudy (anonymous):

what do you mean